Sam drove her car to the mechanic's shop travelling 50 km/h.She left her car to be tuned up and rode home on her bicycle travelling 30 km/h. If the entire trip took 4 hours, how far does Sam live from the mechanic's shop ?
Answer by richwmiller(17219)

a+b=4
50a+30b=x
Not enough info.
x=20a+120 = 20 (a+6)